# Ankle Exoskeletons May Hinder Standing Balance in Simple Models of Older and Younger Adults

### v2

@@ 4.2 Constructing invariant target sets
Recall that controlled invariance is a desirable safety property. To ensure that the BRS is controlled invariant, it suffices to show that the target set is controlled invariant, as explained in Section 2. However, in our case, the target set must be a subset of the safe region defined by constraints (C1) and (C3)-(C5), and obey constraints on the controller (C2). Thus the target sets must be constructed with care.
Appropriate candidates for the target set can be found in the biomechanics literature, where feasible stability is commonly defined as the ability to reach static standing configurations [20]. These are equilibrium configurations where (i) the angular velocity of the body is zero and (ii) the $x$-coordinate of the center of mass ($\text{CoM}_{x}$) is within the base of support (i.e. above the foot). These configurations form a continuum that can be used to construct target sets.
We show this by first analyzing static equilibrium (zero net-torque) states that fulfill condition (i). Let $\tau_{\text{act}}$ denote the total actuation torque about the ankle, as in Figure 2, and let $\tau_{g}=mgl\cos{\theta}$ ) be the gravitational torque. Summing the torques, we see that $\tau_{\text{act}}=\tau_{g}$ meaning that the net actuation torque must remain constant and compensate for gravity. With human torque $\tau_{\text{human}}$ state feedback exoskeleton torque $\tau_{\text{exo}}$, we have that $\tau_{\text{act}}=\tau_{\text{human}}+\tau_{\text{exo}}$. Also recall that in our ‘lifted’ system (3), input $u=\dot{\tau}_{\text{human}}$, is the rate of the human torque development. Therefore we have that at equilibrium, $\tau_{\text{human}}=\tau_{g}-\tau_{\text{exo}}$ and $u=0$.
@@
Our desired static equilibria must also satisfy condition (ii), with the $\text{CoM}_{x}$ above the foot. In angular coordinates, the bounds of the foot can be represented as the maximum forward and backward lean angle, such that the $\text{CoM}_{x}$ is within the foot. We refer to the maximum forward lean as $\theta_{\text{toe}}$ and the backward lean angle as $\theta_{\text{heel}}$. Using the geometric parameters of the foot as shown in Figure 2, we have that $\theta_{\text{heel}}=\cos^{-1}{(\frac{a}{l}})$ and $\theta_{\text{toe}}=\cos^{-1}{(\frac{l_{f}-a}{l}})$. This gives a feasible range of lean angles, $\theta\in[\theta_{\text{heel}},\theta_{\text{toe}}]$.
To model humans realistically, we also account for limits on $\tau_{h}$ shown in constraint (C1). The feasible target states consist of the intersection of the set of states where the constraints on joint angle are satisfied and the set where the constraints on torque production are satisfied. These are shown inside the green square in Figure 3. Note that all static equilibria within this square also satisfy constraints (C3)-(C5). This is because at these states, the system is not accelerating and is at zero velocity - therefore the ground reaction force is equal to the weight of the foot and body, there is no friction force, and the CoP is within the bounds of the foot. Thus we have a continuum of goal positions within the safe region, where the foot maintains contact with the ground without tipping or slipping.
This representation of the target set raises two key concerns. First, because the set consists only of static equilibria, it does not allow for natural postural sway, which is a fundamental characteristic quiet standing in humans. Secondly, one-dimensional target sets consisting of curves or line segments can lead to computational issues when computing the BRS. To generate target sets that are more biologically realistic, we instead use a Control Lyapunov Function (CLF) to construct controlled invariant ellipsoids. CLFs are an extension of Lyapunov functions to systems with control, and are defined as follows [31]:
@@
In other words, a control input exists that stabilizes the system at $x^{*}$. Note that just as the level sets of Lyapunov functions are invariant, the level sets of a CLF are also controlled invariant. The level sets of an appropriately formulated CLF can therefore be used as target sets.
[…]
Because we compute the BRS in two separate forward and backward velocity segments, the target set formulation must encompass all the portions of the foot that we wish to reach with mostly positive and mostly negative velocity only. We therefore make the following observation: If an equilibrium point near the toe can be reached with a positive velocity, then any point posterior to that can be eventually reached with a low enough negative velocity, i.e. by leaning back slowly. This implies that it is sufficient to select one equilibrium point at the toe, rather than having to represent the entire range of static equilibria. A symmetric argument can be made for an equilibrium point at the heel or ankle for points that can be reached with a negative velocity. However, when the system includes a saturating exoskeleton torque due to the device’s motor limits, the dynamics become hybrid. There are three modes corresponding to 1) negative saturation, 2) no saturation, and 3) positive saturation. These modes can be seen in Figure 3. Only two modes appear within the green region. Thus we will select one equilibrium point per mode: a ‘toe’ equilibrium point in mode 1, and an ‘ankle’ equilibrium point in mode 2.
To select the equilibrium point at the toe, we must take into account the functional base of support (FBOS). An individual’s FBOS is defined as the maximum range that they can voluntarily translate their CoP in the anterior (toward the toe) and posterior (toward the heel) directions. In our model, we assume that the angle corresponding to the maximum anterior CoP position is approximately the most anterior point contained within the foot and maximum plantar flexion torque bounds introduced in constraint C1:
@@
The angle $\theta_{\text{toeEq}}$ is largest angle such that the model’s human and exoskeleton maximum torques,$\tau_{\text{human}}^{\text{max pf}}$ and $\tau_{\text{exo}}^{\text{sat}}$, are sufficient to compensate for gravity and maintain a static position. We subtract $0.03$unitrad to allow for displacement due to postural sway. When the model torque bounds correspond to a young individual this point occurs at $\theta_{\text{toe}}$, but in models of older adults, the region of static support may be smaller. Indeed, in an older adult there may be further limiting factors shrinking the FBOS, such as reduced toe flexor strength [32].
The equilibrium point corresponding to the toe is therefore $x_{\text{toeEq}}=[\theta_{\text{toeEq}},0,mgl\cos{(\theta_{\text{toeEq}})-\tau_{\text{exo}}^{\text{sat}}}]$. The ankle equilibrium point is more straightforward. We simply let $\theta_{\text{ankEq}}=\frac{\pi}{2}$. This corresponds to upright standing with no forward lean. The ankle equilibrium is $x_{\text{ankEq}}=[\theta_{\text{ank}},0,\tau_{g}-\tau_{\text{exo}}]$, where $\tau_{\text{exo}}$ is determined by a state feedback law $K_{\text{exo}}(x_{1},x_{2})$. The ankle and toe equilibria for the system with and without exoskeleton assistance are shown in Figure 3, along with the FBOS. Note that in the no exoskeleton case, we simply set $\tau_{\text{exo}}=\tau_{\text{exo}}^{\text{sat}}=0$.
Figure 3: Continua of equilibrium points in the zero-velocity plane for the system with and without an exoskeleton. The shaded green region indicates the bounds of the foot along the $x$-axis, and the human torque bounds on the $y$-axis. The equilibrium point corresponding to the ankle and the points corresponding to the maximum point of static feasible support are marked with blue dots. The bounds in this figure correspond to the older female model.
Figure 4: The top row shows the stabilizable region computed for the young male model (left) and weak older female (right). The baseline regions without the exoskeleton are shown in orange, while the region with exoskeleton assistance is shown in purple. Stabilizable regions when MT and MRTD are reduced independently are shown in the bottom four panels. In the second, the regions are computed with no exoskeleton. The largest sets are baseline, followed by 30% reduction and 50% reduction in each respective quantity. In the bottom row, the exoskeleton is added, and the baseline is outlined in gray.
For a system with a given exoskeleton controller, we linearize the system at each equilibrium point $x_{\text{eq}}\in\{x_{\text{ankleEq}},x_{\text{toeEq}}\}$ such that we have
@@
Where $\tilde{x}=x-x_{\text{eq}}$ is the new state variable representing deviation from $x_{\text{eq}}$. We wish to find some $\tilde{u}=K(\tilde{x})$ that is a closed loop controller guaranteed to stabilize the system about $x_{\text{eq}}$. We can find a CLF of the form $E(\tilde{x})=\tilde{x}^{T}P\tilde{x}$ with a corresponding stabilizing controller by solving a semidefinite program, described in Appendix 8.
The target set centered at $x_{\text{eq}}$ can then be formulated as $\{x\mid(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})\leq c\}$, for some $c\in\mathbb{R}$. However, this CLF is valid for the unconstrained linearized system (14). This means that $c$ should be chosen small enough such that for any $x$ in this ellipsoid, the state and input constraints on the original nonlinear system are satisfied.
To simplify this problem, we first consider only linear constraints. The controller $u=Kx$ should not violate constraint (C2) and the constraint on the human torque production (C1) should also be enforced. Furthermore, the target set should not contain CoM positions anterior or posterior to the foot. As the target set should represent quiet standing, we also add an additional constraint limiting the velocity within the target by an amount close to experimentally observed postural sway [33]. These last two conditions become two additional linear constraints:
$\theta_{\text{heel}}<x_{1}<\theta_{\text{toe}}$ (C6)
$-v_{\text{sway}}<x_{2}<v_{\text{sway}}$ (C7)
where $v_{\text{sway}}$ represents the maximum allowable postural sway within the target set.
The problem of finding the largest $c$ such that the ellipsoid $P$ is maximally contained within the linear constraints is easy to solve. It is a simplified version of the convex problem of finding a maximum volume inscribed ellipsoid within a polyhedron (see for example [34, Chapter 8]) and has a closed form solution:
$c=\min_{i}\frac{(b_{i}-a_{i}^{T}x_{\text{eq}})^{2}}{a_{i}^{T}P^{-1}a_{i}}$ (15)
Here each of the linear constraints is represented as two inequalities of the form $a_{i}x\leq b_{i}$, where $a_{i}\in\mathbb{R}^{1\times 3}$, $b_{i}\in\mathbb{R}$, and $i\in\{1,\dots,8\}$. A complete derivation is in Appendix 8.
For compatibility with the BRS computation tool, we enclose the ellipsoid in a tight axis-aligned bounding box. We check that the vertices are also contained within the constraint polytope. If they are not, we perform a linear search to further scale $c$ such that the constraints are satisfied for the minimum circumscribed axis-aligned bounding box. To account for the nonlinear foot-ground interaction constraints, we check that the box corners do not violate (C3)-(C5). Although this does not provide a provable guarantee of constraint satisfaction, in practice it is a sufficient check due to the relatively small size of the boxes. The target set is then the union of the axis-aligned boxes centered around the ankle and toe equilibria.
@@ 4.3 Evaluating the effect of ankle exoskeleton assistance
To determine how exoskeleton assistance may affect the stability of older adults with differing ankle torque production capabilities, we model five representative individuals. The individuals correspond to the young female (YF), young male (YM), older female (OF), and older male (OM) categories described in [8]. We add the fifth category, weak older female (wOF), by reducing the maximum isometric torque and normalized rate of torque development by half a standard deviation. The respective MT and MRTD bounds of each model are shown in Table 1.
For each model, we construct appropriate target sets as described in the previous section, and then use a state of the art HJB toolbox to compute the backward reachable set in position-velocity-torque space [29]. To form the stabilizable region, we project the resulting BRS into position-velocity space. We first compute a baseline stabilizable region for a young male model without exoskeleton assistance, shown in purple for forward velocities in Figure 1. We validate that for a full strength YM model, the bounds approximately coincide with those computed via the XCoM method shown in the same figure in gray. These bounds overlap wherever the model is close to linear.
@@
To understand the effect of damping, we also compute the SR with a proportional derivative controller whose desired setpoint is static, upright standing:
@@
where $K_{p},K_{d}>0$ represent the PD controller gains. The gains are also taken from [19], where proportional gain $K_{p}=mgl$ is a linearized gravity compensation term, while the derivative gain $K_{d}=0.3\sqrt{ml^{2}K_{p}}$ compensates for gravitational stiffness. Both the GC and PD controller are easily substituted into the dynamics such that $\tau_{\text{exo}}=K(x)$. While we only analyze these two strategies, we emphasize that the methods presented here can be applied to any state-feedback controller.
To get a clearer picture of the independent contributions of reduced MT and MRTD, we also analyze how reduced MRTD and reduced MT affect the SR separately. We compute the SR for the YF model with the MT held constant at the YF’s nominal peak value, while reducing the MRTD in ten percent increments. We repeat this procedure but we reduce the MT while holding the MRTD fixed. We evaluate how deficits in MT and MRTD interact with ankle exoskeleton assistance by computing similarly reduced strength sets with GC exoskeleton assistance added. We compare the area of the reduced strength stabilizable regions with the full strength baseline region, and compute the percent reduction or increase in area.
Table 1: Model parameters for young female (YF), old female (OF), young male (YM). old male (OM), and weak older female (wOF), including MT and MRTD in dorsiflexion (DF) and plantar flexion (PF) directions

### v3

@@ 4.2 Constructing invariant target sets
Recall that controlled invariance is a desirable safety property. To ensure that the BRS is controlled invariant, it suffices to show that the target set is controlled invariant (Section 2). However, in our case, the target set must be a subset of the safe region defined by constraints (C1) and (C3)-(C5), and obey constraints on the controller (C2). Thus the target sets must be constructed with care.
Appropriate candidates for the target set can be found in the biomechanics literature, where feasible stability is commonly defined as the ability to reach static standing configurations [21]. These are equilibrium configurations where (i) the angular velocity of the body is zero and (ii) the $x$-coordinate of the center of mass ($\text{CoM}_{x}$) is within the base of support (i.e. above the foot). These configurations form a continuum, a subset of which can be used to construct target sets.
We first analyze static equilibrium (zero net-torque) states that fulfill condition (i). Let $\tau_{\text{act}}$ denote the total actuation torque about the ankle, as in Figure 2, and let $\tau_{g}=mgl\cos{\theta}$ ) be the gravitational torque. Summing the torques, we see that $\tau_{\text{act}}=\tau_{g}$ meaning that the net actuation torque must remain constant and compensate for gravity. With human torque $\tau_{\text{human}}$ state feedback exoskeleton torque $\tau_{\text{exo}}$, we have that $\tau_{\text{act}}=\tau_{\text{human}}+\tau_{\text{exo}}$. Also recall that in our ‘lifted’ system (3), input $u=\dot{\tau}_{\text{human}}$, is the rate of the human torque development. Therefore, at equilibrium, $\tau_{\text{human}}=\tau_{g}-\tau_{\text{exo}}$ and $u=0$.
@@
Our desired static equilibria must also satisfy condition (ii), with the $\text{CoM}_{x}$ above the foot. In angular coordinates, the bounds of the foot can be represented as the maximum forward and backward lean angle, such that the $\text{CoM}_{x}$ is within the foot. We refer to the maximum forward lean as $\theta_{\text{toe}}$ and the backward lean angle as $\theta_{\text{heel}}$. Using the geometric parameters of the foot as shown in Figure 2, we have that $\theta_{\text{heel}}=\cos^{-1}{(\frac{a}{l}})$ and $\theta_{\text{toe}}=\cos^{-1}{(\frac{l_{f}-a}{l}})$. This gives a range of permissible target lean angles, $\Theta_{\text{range}}=[\theta_{\text{heel}},\theta_{\text{toe}}]$.
Lastly, we account for the limits on $\tau_{\text{human}}$. Feasible equilibria must lie within the intersecton of $\Theta_{\text{range}}$ and constraint (C1) (the green square in Figure 3). Note that static equilibria within this square also satisfy constraints (C3)-(C5). At these states, the system is not accelerating and is at zero velocity - therefore the ground reaction force is equal to the weight of the foot and body, there is no friction force, and the CoP is within the bounds of the foot. Thus a possible target set is the intersection of $\Omega_{eq}$ with the safe region, where the foot maintains contact with the ground without tipping or slipping (Fig. 3, orange/blue curves inside green square).
This representation of the target set raises two key concerns. First, because the set consists only of static equilibria, it does not allow for natural postural sway, which is a fundamental characteristic quiet standing in humans [33]. Secondly, one-dimensional target sets consisting of curves or line segments can lead to computational issues when computing the BRS. To generate target sets that are more biologically realistic, we instead use a Control Lyapunov Function (CLF) to construct controlled invariant ellipsoids. The resulting ellipsoidal target sets encompass a range of realistic sway velocities and are large enough to avoid problems with computational precision.
CLFs are an extension of Lyapunov functions to systems with control, and are defined as follows [34]:
@@
In other words, a control input exists that stabilizes the system at $x^{*}$. Importantly, just as the level sets of Lyapunov functions are invariant, the level sets of a CLF are controlled invariant. CLFs can therefore be used to construct a target set by (i) selecting equilibrium points, (ii) computing CLFs about those points, (iii) determining the appropriate set level and (iv) taking the union of the resulting level sets as the target set.
We start by selecting relevant equilibrium points from (12). When the system includes a saturating exoskeleton torque due to device motor limits, the dynamics become hybrid. There are three modes corresponding to 1) negative saturation, 2) no saturation, and 3) positive saturation. Only two modes appear within the green region, negative saturation and no saturation (Fig. 3). Thus, we choose one point per mode: a ‘toe’ point at negative saturation, and an ‘ankle’ point at no saturation.
To select the equilibrium point at the toe, we must take into account the functional base of support (FBOS). The FBOS is defined as the maximum range that an individual can voluntarily translate their CoP in the anterior (toward the toe) and posterior (toward the heel) directions. In our model, we assume that the angle corresponding to the maximum anterior CoP position is approximately the most anterior point contained within the foot and maximum plantar flexion torque bounds introduced in constraint C1:
@@
The angle $\theta_{\text{toeEq}}$ is the largest angle such that the model’s human and exoskeleton maximum torques,$\tau_{\text{human}}^{\text{max pf}}$ and $\tau_{\text{exo}}^{\text{sat}}$, are sufficient to compensate for gravity and maintain a static position. We subtract $0.03$unitrad to allow for displacement due to postural sway. When the model torque bounds correspond to a young individual this point occurs at $\theta_{\text{toe}}$, but in models of older adults, the region of static support may be smaller. Indeed, in an older adult there may be further limiting factors shrinking the FBOS, such as reduced toe flexor strength [35].
The equilibrium point corresponding to the toe is therefore $x_{\text{toeEq}}=[\theta_{\text{toeEq}},0,mgl\cos{(\theta_{\text{toeEq}})-\tau_{\text{exo}}^{\text{sat}}}]$ (Fig. 3). The ankle equilibrium point is more straightforward, corresponding to upright standing with no lean: $x_{\text{ankEq}}=[\theta_{\text{ank}},0,\tau_{g}-\tau_{\text{exo}}]$. Here $\tau_{\text{exo}}$ is determined by a state feedback law $K_{\text{exo}}(x_{1},x_{2})$. Note that in the no exoskeleton case, $\tau_{\text{exo}}=\tau_{\text{exo}}^{\text{sat}}=0$.
Figure 3: Continua of equilibrium points in the zero-velocity plane for the older female model with and without an exoskeleton. The shaded green region indicates the bounds of the foot along the horizontal axis, and the human torque bounds on the vertical axis. To generate sets representing quiet standing, we construct ellipsoids centered at the ankle and at the point closest to the toe corresponding to maximum feasible static support. These points are marked with blue circles.
Figure 4: Panel (4a) shows the stabilizable region computed for the young male model (top) and weak older female (bottom). The baseline regions without the exoskeleton are shown in orange, while the region with exoskeleton assistance is shown in purple. Stabilizable regions when MT and MRTD are reduced independently are shown in the panels (4b) and (4c). In the top row, stabilizable regions are computed without an exoskeleton. The largest sets are baseline, followed by 30% reduction and 50% reduction in each respective quantity. In the bottom row, the stabilizable regions with exoskeleton assistance are shown in purple, while the reference baseline (full strength, without exoskeleton) is shown in light orange.
We next construct CLFs by linearizing (3) for each equilibrium point $x_{\text{eq}}\in\{x_{\text{ankleEq}},x_{\text{toeEq}}\}$:
@@
where $\tilde{x}=x-x_{\text{eq}}$ is the new state variable representing deviation from $x_{\text{eq}}$. We can find a CLF of the form $E(\tilde{x})=\tilde{x}^{T}P\tilde{x}$ with a corresponding stabilizing human controller $\tilde{u}=K(\tilde{x})$ by solving a semidefinite program, described in Appendix II-A in the supplementary material. Controlled invariant ellipsoids centered at each $x_{\text{eq}}$ can then be formulated as $\{x\mid(x-x_{\text{eq}})^{T}P(x-x_{\text{eq}})\leq c\}$ for some $c\in\mathbb{R}$, with the union of these sets becoming the target set. Note that as each CLF is valid for the unconstrained linearized system (14), each $c$ should be chosen small enough such that the state and input constraints on the original nonlinear system are satisfied for all $x$ in the ellipsoid. Our method for computing appropriately scaled ellipsoids is described in detail in Appendix II-B of the supplementary material.
@@ 4.3 Evaluating the effect of ankle exoskeleton assistance
To determine how exoskeleton assistance may affect the stability of older adults with differing ankle torque production capabilities, we model five representative individuals. The individuals correspond to the young female (YF), young male (YM), older female (OF), and older male (OM) categories described in [8]. We create the fifth category, weak older female (wOF), by reducing the maximum isometric torque and normalized rate of torque development by half a standard deviation of the mean values for the OF in [8]. The respective MT and MRTD bounds of each model are shown in Table 1.
For each model, we construct appropriate quiet standing target sets (section 4.2) and use a state of the art HJB toolbox[31] to compute the backward reachable set in position-velocity-torque space. To form the stabilizable region, we project the resulting BRS into position-velocity space. We first compute a baseline SR for a young male model without exoskeleton assistance (Fig. 1, purple). We validate that for a full strength YM model, the bounds approximately coincide with those computed via the XCoM method (Fig. 1, gray). These bounds overlap wherever the model is close to linear.
@@
To understand the effect of damping, we also compute the SR with a proportional derivative (PD) controller where the desired setpoint is static, upright standing:
@@
where $K_{p},K_{d}>0$ represent the PD controller gains. The gains are also taken from [20], where proportional gain $K_{p}=mgl$ is a linearized gravity compensation term, while the derivative gain $K_{d}=0.3\sqrt{ml^{2}K_{p}}$ compensates for gravitational stiffness. Both the GC and PD controller are easily substituted into the dynamics such that $\tau_{\text{exo}}=K(x)$.
For a high-level understanding of the effect of each control strategy, we compute the total area of the SR without exoskeleton to form a baseline for backward and forward ankle angular velocities. We then determine the total area of the human-exoskeleton SR computed with both GC and PD controller strategies, and calculate the percent change in total area with respect to the baseline. While we analyze two strategies, we emphasize that the methods presented here can be applied to any state-feedback controller.
The three older models in Table 1 combine reductions in MT and MRTD. To get a clearer picture of the independent contributions of reduced MT and MRTD, we also analyze how reduced MT and reduced MRTD affect the SR separately. We compute the SR for the YF model with the MRTD held constant at the YF’s nominal peak value, while reducing the MT in ten percent increments. We repeat this procedure reducing MRTD while holding MT fixed. We evaluate how deficits in MT and MRTD interact with ankle exoskeleton assistance by computing similarly reduced strength sets with GC exoskeleton assistance added. We compute the percent reduction or increase in area of the stabilizable regions for the reduced-strength vs. full-strength models.
Table 1: Model parameters for young female (YF), old female (OF), young male (YM), old male (OM), and weak older female (wOF). Including maximum torque (MT) and maximum rate of torque development (MRTD) in dorsiflexion (DF) and plantar flexion (PF) directions
@@
4.4 Work-energy analysis
We perform a work-energy analysis of a simplified model with and without the GC controller to complement our backward reachability analysis. While less accurate, this allows us to better understand the physical and physiological mechanisms that lead to changes in the SR between the no-exoskeleton and exoskeleton conditions. We assume conservation of energy and use unconstrained and undamped 2-D pendulum dynamics, where $\tau_{h}+\tau_{\text{exo}}=ml^{2}\ddot{\theta}+mgl\cos(\theta)$.
We observe that gravity exerts a negative torque when the CoM is behind the ankle and a positive torque when the CoM is ahead of the ankle. The ‘zero gravity torque’ line at $y=\theta_{\text{ankle}}$ separates the 2d state space into positive and negative gravity torque regions (dashed vertical black line in Fig. 6a). We also define a zero human-torque line:
Definition 5.
Consider first, the no exoskeleton case — The only static standing position where the human does not need to produce a gravity resisting torque is directly above the ankle. Thus, there is a unique ZTL along which total mechanical energy is constant, being equal to the potential energy when the center of mass is directly above the ankle. If the initial energy is $E_{i}=mgl\sin{\theta}+\frac{1}{2}ml^{2}\dot{\theta}^{2}$ and the final energy is $E_{f}=mgl\sin{\theta_{\text{ankle}}}$ ) we can compute the ZTL by letting $E_{i}=E_{f}$, and derive an equation for the system velocity as a function of position:
[…]
This ZTL (Fig. 6a) delineates the direction of the optimal human torque. For a trajectory starting below the ZTL, a positive (dorsiflexion) torque must eventually be applied to avoid a backward fall past the heel. For trajectories starting above the zero-torque line, a negative (plantar flexion) torque is eventually necessary to brake quickly enough to prevent the CoM from overshooting past the toe. Thus, the boundary of the SR that is below the ZTL corresponds to the lower bound before a backward fall occurs, whereas the SR boundary above the ZTL corresponds to forward falls [21]. Based on these fall boundaries, the torque applied by gravity works to prevent falls rather than cause them, in large portions of the state space.
Now we consider how the addition of an exoskeleton with GC control changes the ZTL. Equation (17) is no longer valid once a GC torque is added, as the exoskeleton will apply a torque everywhere except when the gravitational torque is zero. To follow the same trajectory as the original ZTL, i.e. without an exoskeleton, the human would have to exactly counter the exoskeleton torque. Thus, the ZTL changes in the presence of the exoskeleton, and needs to be re-calculated.
Recalling that the exoskeleton torque saturates at specific negative and positive values, we can calculate the angles at which saturations occur. We divide the state space into three regions: with positive saturation, where $\tau_{\text{exo}}=\tau_{\text{sat}}$; no saturation, where $\tau_{\text{exo}}=mgl\cos{\theta}$ ); and negative saturation,where $\tau_{\text{exo}}=-\tau_{\text{sat}}$. The ZTL is computed piecewise for each region.
is the ZTL in the region of positive exoskeleton saturation. The full ZTL with exoskeleton assistance is shown in Figure 6a as a solid purple line.
We can use the work-energy principle to characterize how exoskeleton-induced changes to the ZTL, combined with the human torque bounds, affect stabilizability. The net work of the system is the sum of the work done by the human, exoskeleton, and gravity. Note that the net work done by gravity and the exoskeleton will always have opposite signs due to the controller formulation. For most initial conditions reaching the target set requires a reduction in kinetic energy, i.e. negative work,
[…]
By selecting an appropriate final position $\theta_{f}$, and by assuming that the human produces a constant optimal torque, upper and lower boundaries on the stabilizable region can be computed in a piecewise manner in the position-velocity plane. For example, for an initial angle $\theta_{i}<\theta_{\text{sat}}^{\text{lb}}$, the upper boundary is expressed by the relation
$\tau_{\text{human}}^{\text{max pf}}\cdot(\theta_{\text{sat}}^{\text{lb}}-\theta)+(-mgl)(\sin(\theta_{\text{sat}}^{\text{ub}})-\sin(\theta_{\text{toeEq}}))$
$+mgl\cos(\theta_{\text{toeEq}})(\theta_{\text{sat}}^{\text{ub}}-\theta_{\text{toeEq}}).$